OpenStudy (anonymous):

why is 1/sqrt(3) - 2 equal to -2-sqrt(3)

myininaya (myininaya):

\[\frac{1}{\sqrt{3}-2}*\frac{\sqrt{3}+2}{\sqrt{3}+2}=\frac{\sqrt{3}+2}{3-4}=\frac{\sqrt{3}+2}{-1}\] \[=-1(\sqrt{3}+2)=-\sqrt{3}-2=-2-\sqrt{3}\]